### v1

# The Relational Origins of Rules in Online Communities

Abstract.
Where do rules come from in online communities? While prior studies of online community governance in social computing have sought to characterize rules by their functions within communities and documented practices of rule enforcement, they have largely overlooked rule adoption and change. This study investigates how and why online communities adopt and change their rules. We conducted a grounded theory-based analysis of 40 in-depth interviews with community leaders from subreddits, Fandom wikis, and Fediverse servers, and identified seven processes involved in the adoption of online community rules. Our findings reveal that, beyond regulating behavior and solving functional intra-community problems, rules are also adopted and changed for relational reasons, such as signaling or reinforcing community legitimacy and identity to other communities. While rule change was often prompted by challenges during community growth or decline, change also depended on volunteer leaders’ work capacity, the presence of member feedback mechanisms, and relational dynamics between leaders and members. The findings extend prior theories from social computing and organizational research, illustrating how institutionalist and ecological explanations of the relational origins of rules complement more functional accounts. The results also support design recommendations that integrate the relational aspects of rules and rulemaking to facilitate successful governance across communities’ lifecycles.
copyright: ccjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Preprint under review; May, 2025; posted to https://arxiv.org
1. Introduction
Rules serve many purposes in online communities. Rules codify social norms (Chandrasekharan et al., 2018; Fiesler et al., 2018), signal appropriate behavior (Matias, 2019b), and are subjects of debate in online governance (Gillespie, 2018; Stockinger et al., 2023; Jhaver et al., 2019c; Crawford and Gillespie, 2016). Prior studies in social computing offer insights into extant rules (e.g., Fiesler et al., 2018; Nicholson et al., 2023; Hwang and Shaw, 2022; Keegan and Fiesler, 2017; Reddy and Chandrasekharan, 2023) as well as practices of rule enforcement (e.g., Dosono and Semaan, 2019; Wohn, 2019; Cullen and Kairam, 2022; Matias, 2019b), including the impacts of automated moderation and enforcement tools (e.g., Chandrasekharan et al., 2019; TeBlunthuis et al., 2021; Halfaker and Geiger, 2020). However, most studies of rules take for granted that rules must also be designed and adopted, yet we know little about the origins of rules in online communities. Understanding the origins of rules is important because it reveals a design space of rulemaking and community governance.
There are two broad approaches to understanding rulemaking in the social computing literature. First, a top-down “social engineering” perspective (Kraut and Resnick, 2012) aims to identify and evaluate generalizable principles that platforms can implement to regulate behavior and guide effective content moderation (Chandrasekharan et al., 2019, 2022; Matias, 2019b; Jhaver et al., 2021; Seering et al., 2017; Horta Ribeiro et al., 2021). An alternative “institutional analysis and development” (IAD) or ”institutional design” perspective (Frey et al., 2019b) emphasizes the bottom-up and localized development of rules, rulemaking norms, and enforcement practices within communities (e.g., Schneider et al., 2021). Despite their divergent orientations, both approaches assume that rules are primarily functional—i.e., created to solve problems located in the community adopting a rule.
Our alternative approach begins from the observation that many online communities share similar rules (Reddy and Chandrasekharan, 2023; Fiesler et al., 2018; Nicholson et al., 2023; Hwang and Shaw, 2022). A functional account of rulemaking would attribute these patterns to common problems communities face. However, research in organizational sociology has rejected such explanations of rule similarity in firms since the 1970s. This work—including “new institutionalism” (e.g., Meyer and Rowan, 1977; DiMaggio and Powell, 1983; Tolbert and Zucker, 1983) as well as “organizational ecology” (Hannan and Freeman, 1984; Carroll, 1985)—has argued that similarities in organizational structures are best explained in reference to external social pressures. We apply these insights to the crafting of the rules in online communities.
To understand the origins of rules, we conducted 40 in-depth, open-ended interviews with leaders of communities from three different sociotechnical systems: subreddit communities on Reddit, wikis on Fandom, and servers in the Fediverse. Interviewees were recruited from clusters of communities with similar rules in each setting. We report findings based on a comparative, grounded theory analysis (Charmaz, 2015) that draws on the earlier work in organizational sociology as sensitizing theories.
We find that rules originate and change for a mix of functional and relational purposes that go beyond merely providing the basis for enforcement decisions or conveying community goals and norms. Rules and rulemaking play important, underappreciated roles as communities navigate relationships with internal and external constituencies and constraints. In particular, we find that community leaders use rules to draw distinctions between communities, manage issues of community growth as well as decline, and legitimize the community to external peers and platforms.
Our paper makes several contributions to social computing theory focused on governance. First, our account of the origins of rules emphasizes how rules mediate online communities’ relationships to other communities, platform rules and affordances, and audiences of potential and current members, all of which constrain rulemaking. This relational aspect of rulemaking extends prior functionalist accounts and expands understanding of the design space of digital institutions and community self-governance (Frey et al., 2019a; Zhang et al., 2020; Frey and Schneider, 2021; Jhaver et al., 2021; Schneider, 2022). Additionally, we synthesize concepts from new institutionalism and organizational ecology to show how rules play a relational role in community ecosystems by communicating niche identity and legitimacy. Our work also makes empirical contributions with practical importance by filling a critical qualitative gap on the origins and practices of rule creation and change in online communities and highlight opportunities for rule adoption strategies across community lifecycle phases. We also provide comparative empirical evidence across three different systems to show how technological design can influence rule adoption and change.
2. Background
Rules are central to governance in online communities. Yet, when research in social computing has considered rules, it typically takes for granted how rules are initially adopted and how they evolve. Our paper seeks to build new theory about why and how rules in online communities are created and changed. Although our research design is inductive and our findings emerge from our data, there are two lines of existing scholarship that frame our study. The first line includes two major paradigms in prior social computing work on online community governance: the “social engineering” and “institutional design” perspectives. This scholarship provides the rationale for our study as well as its design. The second line of research comes from the “new institutionalist” and “organizational ecology” traditions in sociology. These literatures help us interpret and synthesize the themes that emerged from our analysis.
2.1. Governance and Rules in Online Communities
Governance is the social practices and processes that shape behavior within and between institutions (Bednar, 2023). As part of governance, rules are the explicit “norms, regulations, and expectations that regulate the behavior of individuals and interactions among them.” (March et al., 2000, p. 5). Social computing research has explored rules and governance in online communities—the virtual spaces where individuals engage around shared interests or goals through socially regulated interaction on a social media platform (Jones, 1997; Preece, 2000; Leimeister et al., 2004). [^1] For our purposes, a “social media platform” is a website that affords the creation of communities around user-generated content and interactions (Carr and Hayes, 2015) This study responds to two influential perspectives in social computing on rules in online communities.
The first is the top-down “social engineering” approach that centers the role of “designers,” and is exemplified by the work in Building Successful Online Communities (Kraut and Resnick, 2012). This research paradigm seeks to identify generalizable principles by which platform designers can shape behavior in ways that correspond to their goals. For example, Kraut and Resnick (2012)’s chapter on norms and rules lists a series of “design principles” that offer specific guidance for how a community’s designer can set injunctive norms (i.e., rules) to shape participants’ behavior. Other studies test how rule placement can increase salience and compliance (Matias, 2019b), evaluate how post-enforcement rule explanations impact subsequent user behavior (Jhaver et al., 2019c), and investigate the effectiveness of different rule enforcement mechanisms (Horta Ribeiro et al., 2021; Chandrasekharan et al., 2017). A significant line of research also focuses on describing existing governance approaches and developing new ones (Geiger, 2014; Seering et al., 2017; Chandrasekharan et al., 2019; Jhaver et al., 2019a, b; Zhang et al., 2020).
The second perspective critiques the first by focusing on the role of end-users who create their own systems of governance, even as they may lack full control (Kiene et al., 2019; Seering et al., 2019; Frey et al., 2019b). This “institutional design” approach draws on the Institutional Analysis and Development (IAD) framework created to study common pool resource governance by Ostrom (1990) and colleagues. It highlights a variety of backstage actions and decisions, as well as the complex internal organization of online communities (Butler et al., 2008; Frey and Sumner, 2019) and embraces the diversity of problems, resources, and goals across communities (Keegan and Fiesler, 2017; Fiesler et al., 2018; Nicholson et al., 2023; Reddy and Chandrasekharan, 2023; Hwang and Shaw, 2022). As such, the institutional design perspective favors localized, community-specific systems and solutions over top-down social engineering.
In the social engineering approach, changes to the rules are a reflection of the designer’s changing desires, making rule change theoretically equivalent to rule adoption. Because the institutional design approach tends to be focused on durable governance over time, rule change is a much more important consideration. Ostrom (2005) points to the importance of a “constitutional” layer that dictates the rules by which rules can be changed. In their study of social computing systems, Frey et al. (2019b) invoke the need for a constitutional layer to argue that effective governance depends on institutions that can adapt to changing circumstances.
Existing work on online communities has often approached the topic of rule change from the perspective of contestation. Scholars note how moments of rule enforcement can become opportunities to contest the meaning and understanding of rules (de Laat, 2012; Crawford and Gillespie, 2016; Matei and Dobrescu, 2010; Beschastnikh et al., 2008; Kriplean et al., 2007). That said, the little empirical evidence we have on the subject suggests that written rules in online communities change infrequently. For example, Hwang and Shaw (2022) finds that rules on the five largest language editions of Wikipedia are increasingly less edited even as they continue to be discussed. Evidence from the broader social computing literature offers some explanations for why rule changes may not be frequent online. Online communities are numerous and relatively cheap to reproduce (Benkler, 2006). This might make opportunities for rule change less likely if members can simply leave and join another community that already has the rules that they want. Hirschman (1970) suggests that the presence of alternatives can influence whether leaders ignore problems or enact change. Moreover, scholars have noted that opportunities for “voice” in online communities are limited (Schneider, 2022; Frey and Schneider, 2021; Shaw and Hill, 2014).
A key assumption of both the social engineering and institutional design perspectives is that rules are functional and that rule selection and change can be understood as the output of rational processes: choices to adopt or change rules are determined by specific problems within specific communities. From a functional standpoint, efforts to understand governance in both perspectives focused on evaluating whether rules or rulemaking successfully mitigate internal problems. For example, research on content moderation (Grimmelmann, 2017) across the two perspectives is centrally concerned with the effectiveness (Chandrasekharan et al., 2019) and consequences (Jhaver et al., 2019a) of rule enforcement practices. Both perspectives treat rules as emerging to address these internal problems, just from different sources. To the extent that communities wind up with similar rules, it is because they have similar problems that benefit from similar solutions. The sources of rules have largely remained unexamined in both streams of research.
2.2. Sociological Perspectives on Organizational Similarity
Research in organizational sociology plays two important roles in our study. First, it provides a set of arguments for why organizational rule sets might look similar—what Charmaz (2015) calls a “sensitizing theory,” that might explain why so many different online communities share rules. Second, although we did not embark on our study with the expectation that these theories would figure heavily in our description of rule adoption and change, we found that they provide an excellent theoretical vocabulary for describing our empirical findings. We introduce key concepts from this literature here that we rely on in our Findings and Discussion.
The most important body of scholarship in organizational sociology we rely on is known as the new institutionalism. It was first articulated by Meyer and Rowan (1977) and DiMaggio and Powell (1983). Confusingly, there are several scholarly traditions in economics, political science, and sociology referred to as “institutionalism” and even “new institutionalism” (Powell and Dimaggio, 1991). These traditions have distinct emphases and often define institutions differently. For example, the institutional design perspective described in §2.1 draws from North (1990), Ostrom (1990), and others associated with what is sometimes called “new institutional economics.” This approach defines institutions as “rules of the game” and seeks to increase attention to the way written and unwritten rules shape and constrain the behavior of individuals and organizations.
New institutionalism in sociology is a distinct and more conceptually narrow line of work that argues that organizations’ rules and routines do not reflect the rational solution to problems faced by organizations, something taken for granted in much of the work on rules in economics. They are instead based on “myths” (Meyer and Rowan, 1977) and are the results of isomorphic pressures that cause organizations to adopt similar structures, rules, and routines as a way to earn social legitimacy (DiMaggio and Powell, 1983). DiMaggio and Powell (1983) argue that organizations will copy routines from other successful groups (mimetic isomorphism), adopt them because they reflect expectations from groups like professions (normative isomorphism), or choose them because outsiders like investors or governments require them (coercive isomorphism). Although specific sources of isomorphic pressure differ across organizations and environments, this line of scholarship provides an obvious entry point to approach our observation of similarity in online community rules. After all, online communities on a given platform face similar constraints because they are on the same social computing system (Caplan and boyd, 2018), are subject to similar external policies (Tosch et al., 2024), and rely on similar tools (Hwang et al., 2024).
A distinct tradition within organizational ecology treats organizational behavior as contingent on environmental factors that can promote either variation or conformity in organizational structure (Cyert and March, 1963; Hannan and Freeman, 1977). This “organizational ecology” perspective seeks to explain how organizations may differentiate themselves to avoid directly competing with one another and to meet diverse and specialized needs of audiences or markets (Carroll and Swaminathan, 2000). While the new institutionalists traditionally focused on explaining organizational similarity (i.e., why are organizations so similar?), ecologists more typically focused on variation (i.e., why are organizations so different?). The organizational ecology approach has been echoed by social computing scholars who note that communities can range immensely in their topics, scales, goals, and norms (Chandrasekharan et al., 2018; Weld et al., 2021; TeBlunthuis and Hill, 2022; Zhu et al., 2014, 2013; Hwang and Foote, 2021). Drawing explicitly from organizational ecology, TeBlunthuis et al. (2022) observes how normative differences between communities with similar topics enable communities to be mutualistic rather than competitive with one another. Related work has mapped out “ecologies” of the rules and values that would underpin such normative distinctions (Chandrasekharan et al., 2018; Weld et al., 2021; Fiesler et al., 2018; Nicholson et al., 2023; Hecht and Gergle, 2010). Although researchers have discussed deliberative processes and bureaucratization around rules within specific communities (Black et al., 2008; Butler et al., 2008), we do not know of work that seeks to explain how ecologies of community rules form in online contexts.
An important distinction between the new institutionism and ecology traditions is in the way they approach change. For new institutionalists, organizational change is a central object of study. For example, a new law or regulation may pressure a firm to change its rules to comply with new legal requirements. Similarly, a nonprofit might change its rules to qualify for funding or to conform to published best practices. A key implication from the new institutionalist perspective is that organizations will face similar pressures that will make them structurally similar over time. Conversely, organizational ecologists rely heavily on the concept of structural inertia, which assumes that organizational change is difficult or even impossible (Hannan and Freeman, 1977; Young, 1988). If organizations appear similar, it is because they face similar selection pressures, which have caused dissimilar organizations to fail, or because there are positive returns to legitimacy that stem from similar organizations in an environment. That said, some organizational ecologists have argued that certain features of environments may lead to more organizational change (Hannan and Freeman, 1984).
Although the new institutionalist and organizational ecology perspectives are quite different in their assumptions and predictions, both perspectives share a key sociological insight that is largely absent from most scholarship on governance in social computing: the origins of organizational forms—including rules—lie not only in function but also in the relationships between the organizations and their environments. According to both perspectives, we cannot fully understand communities by looking within communities or by focusing on the problems that the text of the rules purports to solve. In both perspectives, rules are deeply social and communicative.
3. Empirical Settings: Subreddits, Fandom wikis, and Fediverse servers
We study online communities drawn from three settings—all the site of prior social computing work on governance and rules. Although most social computing studies focus on a single sociotechnical system at a time, we chose multiple systems to support a more general, comparative account of rulemaking. Looking across settings allows us to consider how the design of systems might impact rulemaking processes. The three systems we selected each include distinct populations of online communities that are created and managed by community members (users) rather than platform owners: discussion communities (subreddits), peer production communities (Fandom wikis), and federated microblogging communities (Fediverse servers). Below we offer a brief introduction to each system and the ways each are governed.
3.1. Subreddits
Subreddits are user created forum spaces on the networked community platform Reddit.com. A Reddit user creates a subreddit by giving it a name and a topic. Other users may contribute posts (threads to start off a discussion) or comments (to discuss those posts). Subreddits have been created to support a wide variety of topics, from posting pictures of cute animals to discussing the implications of AI. Reddit users can search for subreddit communities and join by subscribing to a feed that shows the user posts from subreddits they follow.
Governance of subreddits is primarily enacted by the creator of the subreddit and other volunteer content moderators they appoint (Seering et al., 2019). Reddit provides community leaders with a suite of tools, including automated rule enforcement with the AutoModerator bot (Jhaver et al., 2019b), membership management, and a member communication system called Mod Mail. Subreddit leaders often prominently display their rules on the side of the subreddits’ main page and use built-in wiki features to elaborate on community guidelines and FAQ’s. Subreddit rules usually have to do with community topic or behavioral expectations (Fiesler et al., 2018). Studies in social computing have mapped out the kinds of rules that are adopted across subreddits (Chandrasekharan et al., 2018; Fiesler et al., 2018) and quantitative factors for rule change (Reddy and Chandrasekharan, 2023).
3.2. Fandom wikis
Wikis are collaboration-based websites where users peer produce public information repositories, like an encyclopedia, by directly creating, editing, and organizing content on the website’s pages. Fandom is a platform that affords the creation and management of wikis, often made up of members of a fandom who share a strong interest in documenting information about a specific work of fiction or cultural artifact (e.g., TV show, book series, or video game). Interactions primarily occur on the “Talk Pages” of articles in a Fandom wiki, where members discuss article content. Some Fandom wikis choose to activate a feature in the platform that provides the community with a simple discussion forum.
Governance in Fandom wikis is more structured and often results in hierarchical roles. As the goal of wikis is to produce relevant, high-quality articles documenting their fandoms, leaders frequently develop rules regulating article content. Users who create a Fandom wiki become its de facto leader but frequently appoint other community members into leadership positions. Social computing scholars have documented how wikis engage in ongoing deliberation about how to change their communities’ rules (Keegan and Fiesler, 2017), and shown that as they grow and mature they become more bureaucratic (by adding more rules) (Butler et al., 2008) and oligarchic and closed to newcomers (Shaw and Hill, 2014; Halfaker et al., 2013; TeBlunthuis et al., 2018).
3.3. Fediverse servers
The Fediverse is an open, decentralized network of servers (also called “instances”) that run the ActivityPub protocol. Users join the Fediverse by registering an account on a specific server of their choice. The server enables its members to share and receive content from other servers, allowing users to interact with fellow server members and the broader network through comments, reposts (“boosts”), likes, and more. We focused on platforms using the Mastodon and Pleroma software, which both offer a microblogging experience similar to systems like Twitter. Servers we observed were generally organized around a shared interest or identity, but members were often not required to post content related to that theme.
Envisioning decentralized social media ecosystems without centralized owners, Fediverse servers using Mastodon and Pleroma are governed exclusively by the servers’ creators, who are referred to as “admins.” Admins control membership through open, closed, or application-based user registration, moderate content through admin panels and back-end access to data, and can block other servers to prevent direct interactions between users of their server and another (known as “defederation”). Rules on Fediverse servers are usually about speech, focused on civility and legality. Although scholarly attention to Fediverse is relatively nascent, researchers have examined migrations to decentralized platforms (Dunbar-Hester, 2024; Cava et al., 2023; He et al., 2023; Nunes, 2023; Jeong et al., 2024), how platforms conceptualize decentralization (Zulli et al., 2020; Gehl and Zulli, 2023; Raman et al., 2019), community governance challenges (Anaobi et al., 2023; La Cava et al., 2024; Bono et al., 2024; Hwang et al., 2025; Zhang et al., 2024), coordination against hate speech (Caelin, 2022), rules on Mastodon platforms (Nicholson et al., 2023), and the effects of defederation on activity (Colglazier et al., 2024).
4. Methods
This work was a collaboration involving professors and graduate students from [Anonymized for Review]. In the rest of this section, we outline the steps taken by our team to collect and analyze data.
4.1. Community Sampling and Participant Recruitment
To investigate the origins of rules, we sought a sample of communities that varied in terms of goals and setting. We focused on three different settings introduced in the previous section—subreddits, Fandom wikis, and Fediverse servers. Between 2020 and 2021, we began by building datasets of community rules from each of these types of sociotechnical systems. Because we were initially motivated by observations of rule similarity across communities (Hwang and Shaw, 2022; Fiesler et al., 2018; Reddy and Chandrasekharan, 2023; Nicholson et al., 2023), we computationally clustered communities within each system by their rule language. We then screened clusters to identify a diverse, theoretically relevant sample of active communities to target for interviews. A more detailed description of this process can be found in Appendix Community Sampling Methods.
After receiving IRB approval for human subjects research from [Anonymized for Review], we contacted moderation teams and admins from the identified communities using direct messages on their respective platforms. In total, we recruited 40 community leaders across 34 communities: 9 from subreddits, 14 from Fandom wikis, and 17 from Fediverse servers. A table describing the communities we recruited from can be found in Appendix A.
4.2. Interviews
We conducted semi-structured interviews via video calls between June and September 2022. All participants were either volunteer moderators or administrators who had familiarity with their community’s rule history and were above the age of 18. Interviews ranged from 30 to 120 minutes in length, averaging 60 minutes. A combination of trained undergraduate research assistants and the paper authors conducted the interviews following an interview protocol that asked participants to describe or tell us stories about their interactions with rules in their communities (see Appendix C). We asked questions about the origins of their rules, how they were enforced, when they changed and why, and about related challenges in governing online communities.
4.3. Analytical Process
Our analytical process was based on Charmaz’s approach to grounded theory (Charmaz, 2015), thematic analysis (Braun and Clarke, 2006), and collaborative affinity mapping. Small teams of trained research assistants led by four authors conducted line-by-line coding of interview transcriptions using the Taguette software, inductively coding text to break down participant’s experiences into actions and feelings through gerunds (e.g., “relaxing the rules” or “feeling a sense of belonging”). Together, we developed over 4,500 open initial codes. Discussions across the authorship team led to a rough list of 20 themes these codes covered. Guided by this list, three authors then conducted a round of focused coding, a step which merges similar initial codes together and elevates them with broader conceptual labels. This led to 337 unique focused codes.
Three authors then collaboratively engaged in an in-depth affinity mapping exercise with the 337 focused codes to visualize patterns, themes, and processes via a digital whiteboard. Rounds of memo-writing and weekly meetings for discussion and iteration complemented this process. The three authors reached agreement on the most significant contributions to bring forward in this paper: rule adoption and rule change. Finally, the first author engaged in a round of theoretical sampling by using the list of 337 focused codes as a codebook to apply to all 40 interview transcripts. We did this to identify supporting evidence across all three empirical settings and to analyze differences when patterns varied between them. The first author distilled a core set of processes and beliefs around rule adoption and change from this analysis, which was iterated upon with feedback and input from the other authors.
5. Findings
From our interviews, we identified seven distinct processes involved in the production of rules in online communities. The first three processes primarily relate to the adoption of an initial rule set. The latter four processes focus on changing a community’s rule set.
5.1. Searching for Rule Templates
Few leaders we spoke to imagined a rule set during community creation. Instead, most leaders described the action of searching for rule templates in other online communities. For example, FA14, the leader of a Fandom wiki related to Lego, the plastic construction toy, told us that they “are always inspired by other communities” and “always take a template of rules” from other communities. Some leaders described feelings of not being “actually experienced enough and knowledgeable enough” (FV12) to write rules themselves. FV10, a Fediverse leader for a regional community in the U.S., told us how “a big chunk” of their rules were made at community creation, saying that they “looked to other examples of rules” across Fediverse communities. They referenced their legal background to explain choosing templates over drafting rules themselves:
“As a lawyer, it’s pretty typical to look at other example documents if you’re writing a contract or something. You look at other examples of how people have done it…And that’s kind of how our rules developed. You see what challenges other people have encountered, and then you can assess whether or not that’s a challenge that you think that you need to be prepared for.”
FV10 showed that using existing rule templates from related communities helps leaders anticipate potential challenges and craft suitable policies for their needs.
5.1.1. Regulating Content and Behavior
In the process for searching for rule templates, leaders primarily cited two motivations: regulating content and regulating behavior. For subreddit leaders like R1, rules regulating content were important for facilitating meaningful discussions. R1 explained why their Korean drama fan community’s rules about mandating substantial discussion posts were important:
“You can’t just write one or two sentences…If you’re going to do a discussion topic, you can’t just say like, oh, ‘What’s your favorite romance?’ You have to give us an answer to that already to give us a jumping off point for everyone else to join in on the discussion.”
For R1’s community, putting effort into posts beyond just a sentence or two was crucial for stimulating deeper conversation and promoting the community’s values.
Leaders on subreddits often implemented rules on post formatting and content. For example, some communities had strict rules on advertising, such as R4’s cybersecurity discussion community. Others had regulations on content they called “misinformation,” such as R5’s regional community, which had rules adopted to prevent the spread of conspiracies related to the COVID-19 pandemic. We found that Fediverse servers rarely required specific kinds of content in their rules. However, they did restrict kinds of content that are inappropriate for professional work settings (NSFW content), hateful content, and advertising.
Fandom leaders frequently saw their rules as setting the standards for contributions to wikis. For FA6, a leader of a Fandom wiki for a popular multiplayer video game, rules governed information presentation and standards for accuracy. This was typical of rules across the Fandom wikis, almost all of which have content requirements and restrictions on formatting, article relevance, and accuracy.
Behavioral regulation was another important justification for rule adoption across all three settings. For subreddit leaders, this was expressed by a desire to ensure discussion spaces remained civil and safe. R9, a leader of a subreddit that organizes trades of living plants, summarized the sentiment that we heard from other subreddit leaders: “I think the biggest purpose of the rules is to keep everybody safe—to make sure nobody’s getting scammed…to make it so the community doesn’t become a toxic place.”
Although Fandom wikis place less emphasis on facilitating discussion, leaders still adopted rules to promote civility in spaces such as article “Talk Pages” and discussion forums, and in affiliated Discord servers that some Fandom communities used for synchronous chat. FA11 summarized the justification for behavioral rules we heard from other Fandom wiki leaders: “They’re there for safety, and they’re also there to make people feel like they’re included and they belong.”
Finally, behavioral regulations were especially important for many Fediverse leaders, many of whom had migrated from Twitter to the Fediverse due to concerns about behavior. FV7 created their community on Mastodon because Twitter was “too toxic.” FV8 felt there were no rules in place on Twitter to protect their community, citing “a lot of abuse and harassment” that their members experienced. Many leaders, like FV16, adopted rules to protect their members from specific kinds of harassment and hate speech, such as racism or transphobia. FV10 explained how rules restricting such behaviors helped the community advertise to potential newcomers the community’s “vibe” and helped “creat[e] expectations for people who might want to join.” These leaders chose initial rules for functional reasons: to establish expectations of acceptable behavior and content and in turn to structure participation to achieve their communities’ goals.
5.2. Copying Other Communities’ Rules
Leaders gave various accounts of how they adopted rules for their communities after searching and identifying potential templates. They often mentioned two distinct next steps: copying other communities’ rules and customizing the adopted rule set5.3). One of the original questions motivating our study was “Why do so many online communities have similar rules?” In the previous section, we showed how leaders would seek out other communities’ rule sets to use as starting templates. For many, the next step was to copy. This pattern was most evident in our interviews with members of the Fediverse. FV10 explained this step as “copying the good stuff” after they “look[ed] at other examples of how people have done it.” Leaders would copy and paste rules they felt were a good match for their community and then later “sort of build on it” (FV10) with further customization to their communities’ unique needs. That said, many leaders from the Fediverse would copy the entire rule set of another community and never customize or change it. In explaining these actions, these leaders emphasized a relational function of rules with the goal of establishing external legitimacy among other communities in their decentralized networks.
5.2.1. Establishing External Legitimacy
A key motivation behind the process of copying other communities’ rules was the desire to project legitimate governance (i.e., the presence of responsible content moderation) and communicate alignment with established norms across an ecosystem of online communities. Fediverse leaders have the power to block (“defederate”) communities that threaten their well-being and goals. These same leaders felt their choices were being observed and judged by other admins of other Fediverse servers that might need to consider whether to defederate at some point. FV10 said this adds “an important function of rules” because it “sends signals to other admins that they know what we are moderating.” This external scrutiny from other admins motivated Fediverse leaders to adopt rules because these rules would signal accountable governance and content moderation. In this way, rules help establish external legitimacy for communities in decentralized networks.
Many Fediverse leaders pointed to mastodon.social’s rule set as a benchmark of legitimate community governance. mastodon.social was one of the first communities on the Fediverse and was created by Eugen Rochko. Rochko is the developer of Mastodon, one of the most widely used platform for hosting Fediverse communities. Fediverse leaders often looked to communities like mastodon.social because it’s “the largest instance” and because it’s “hosted by Eugen” (FV5). FV5 further explained, “You know, if you’re gonna copy rules about a new community, you might as well go right from the source. So I copied most or all of the rules.” The “source” in this case was mastodon.social. FV2 came to a similar conclusion:
“I went to other instances. I looked at what they had in terms of rules, and the rules from the mastodon.social instance made sense to me. And I also saw a lot of other instances referring to those rules, which kind of created this network effect.”
The observation that so many other communities on the Fediverse also referenced mastodon.social’s rule set was an important reason that FV2 adopted that particular rule set. From what we can infer from FV2, the widespread adoption of mastodon.social’s rules created a kind of self-reinforcing dynamic—communities that mirror those rule sets signal their alignment with the standards and norms of a reputable, established community.
In addition to borrowing from mastodon.social, several Fediverse leaders integrated the Fediverse-Friendly Moderation Covenant (FFMC) into their rule sets as a pledge to adhere to shared, cooperative standards of governance with other Fediverse admins. As FV1 noted, adopting the FFMC not only signals a commitment to cooperative governance between servers but also aligns a server with a “bubble of known instances,” establishing the admin as “a stand up citizen, a stand up member of the Mastodon community.” These alignments, whether they are achieved by adopting mastodon.social’s rule set or pledging to the FFMC, builds trust among peer admins and lends emerging communities on the Fediverse—who must federate with other communities to share content—the external legitimacy they believe is needed.
Subreddit and Fandom wiki leaders did not mention copying other communities rules to signal legitimacy in the way that Fediverse leaders did. That said, some Fandom Wiki leaders like FA3 sought out rules from “other big wikis” to see how others addressed similar issues and used their rules as starting points. For some specific rules related to behavior, subreddit leaders like R4 and R5 mentioned copying guidelines from the “Reddiquette,” which is an informal set of behavioral guidelines written by other Reddit leaders similar to the FFMC.
5.3. Customizing the Adopted Rule Set
Customizing the adopted rule set was a natural next step for many leaders. After all, the communities from which they found their initial rule sets often differ in their specific goals. Several interviewees from Fandom and Reddit offered insight into this process. FA1, a leader of a Fandom wiki documenting aesthetic fashion and cultural trends, explained that “it would be nice” if every other Fandom wiki had the same rules, but that rules have to be modified to each community’s unique needs: “every community is different.” Leaders often searched for specific kinds of rules from different communities to solve their problems. FA4, the leader of a Fandom wiki cataloging the lore of a popular comic book franchise, described this process in the context of creating a rule on image usage by customizing the image policy from another wiki:
“We didn’t just copy the rules. We saw exactly what those other communities were asking of their users—and we saw what worked for us, what didn’t, and what the other sites were not asking. And then we adjusted to our specific community.”
We note that changes were not driven purely based on differences in community topic. Leaders also described trying to improve the language or normative effect of a rule. For example, R4 based their rules on other communities they had led and made them “better to a certain degree.” Sometimes leaders described these edits to other rules as minimal and changing them “a little bit.” (FV12) Other times, the process of rewriting the rules could be “agonizing” (FV11) as leaders struggled to make their rules sound “well written.”
5.3.1. Reinforcing Community Identity
The process of customizing the adopted rule set revealed another motivation behind why rules exist: reinforcing community identity. While rules that structure community identity fulfill a functional purpose by restricting contributions that don’t achieve community goals, they also serve a relational purpose by defining the boundary between what the community is and what the community is not—setting a community apart from others that overlap in membership or topic. This was especially important for subreddits and Fandom wikis, where communities with similar topics attract the same audiences who may arrive with misaligned expectations about core community goals. For example, R4 adopted a rule requiring that contributions “Must be relevant to cybersecurity professionals” because newcomers would arrive asking for cybersecurity troubleshooting. R4 explained the rationale for this rule: “we’re trying to create a community for technical professionals.” Rules like these helped to make explicit the community’s intended audience, reinforcing the identity of their community. R4 further reinforced r/cybersecurity’s identity by creating an offshoot subreddit specifically designed for aggregating questions about cybersecurity help and directed newcomers seeking help to post there, linking to it in r/cybersecurity’s sidebar.
In another example, R2 manages two subreddits: both involved on Q&A style discussions, but they differ in the level of formality and seriousness. R2 explained that the rules in r/explainlikeimfive have to be strict and structure a “rigorous environment” while the rules in r/nostupidquestions allow for more freedom in what can be asked and how it can be answered, noting that the rules help keep that community’s discussions “as open as humanly possible.” R2 said “We [at r/nostupidquestions] want a place that’s not r/explainlikeimfive.” By establishing these distinct expectations with their rules, both communities thrived despite overlapping a similar niche space. The rules helped distinguish the communities by the style of Q&A format: one for formal and serious discussions and one that’s completely informal and casual. The tailored rules communicate to both members and newcomers not only what the community is, but also what it is not.
Rules also helped reinforce identity in Fandom communities. For example, FA9, a leader for the DC Wiki (a fandom for the comic book franchise), compared their rules to the rules of the Marvel Wiki (a fandom for a different comic book franchise) to explain how their community was unique. Both wikis utilized a Fandom feature, which adds a forum for members to have discussions about their wiki’s content. That said, FA9 noted that in the DC Wiki, rules there promote “meaningful discussions” by restricting content related to fan fiction, while the Marvel Wiki has more “permissive” rules, allowing its members to discuss a variety of things, including fan fiction. FA9 lamented the disruption from newcomers expecting rules like those in the Marvel Wiki’s community, telling us that the DC Wiki is “not the place” for that kind of content. Like R2, FA9 used rules to explain what their community is and isn’t about: members are supposed to stick closely to discussing content published in the comic book franchise and nothing else, including other media sources and fan fiction.
Identity construction from rules was less explicit coming from Fediverse leaders. While some communities portrayed an identity (e.g., a Mastodon server for people in the Bay Area of California or a community for amateur radio enthusiasts), their rules usually did not limit discussions based on their community’s identity. However, some Fediverse leaders repeatedly mentioned that rules restricting discrimination and hate speech helped to signal a virtue of intolerance to intolerance. In other words, some Fediverse leaders saw the absence of such rules in other communities as an indicator that they allow hate speech to flourish. There were called “free speech” servers because it was inferred that content moderation was virtually absent. FV13 said leaders of free speech communities “don’t care about moderation–[they let you] say what you like.” FV8 told us that they adopted their strict rules on discrimination and hate speech to “signpost” to outsiders that their community is not a place for “free speech warriors.” In one instance, FV14 said their rules essentially help people self-select out because registration requires them to read their anti-discrimination rules.
5.4. Encountering Problems that Challenge the Rules
When asked why the rules were changed, leaders tended to give functional accounts by describing how various problems would raise questions of whether adopted rules sufficed to achieve community goals. However, not all communities experienced such problems or felt the need to change rules at all—particularly those closed to new members or low in activity. For FV17, an absence of problems in their community resulted from the tight-knit and homophilous nature of the group who “kind of know each other already”. Notably, on the Fediverse, leaders can close account registrations on their server and thus control who may become a member. For example, FV10’s server “pretty much never had open registration” and thus never had to change rules. Similarly, relatively small or new communities, like FA6’s wiki community, rarely saw problems with rule breaking because “the number of edits was in the very low volume.” When membership and activity are low, there are less opportunities for problems to emerge and threaten community goals. But even if a community has not experienced problems, this does not mean it never will. Our interviewees discussed how a community might encounter problems—and then change its rules in response to them—in varying ways over the community’s lifecycle, particularly as communities grew or declined.
5.4.1. Problems from Growth
As communities grew in membership and activity, leaders increasingly had to consider whether the community is still able to achieve its goals under the old rule set. In particular, as communities grow and the number of people attempting to contribute increases, content-related issues tend to emerge and occasion leaders to reevaluate and reinforce their communities’ goals by making rules controlling quality and identity.
For example, FA1’s community goals are stated “as a place to know information about online and offline trends, communities, fashions, etc.,”, and they had to adapt their rules to maintain the quality of their wiki in light of this goal. When an influx of new people created articles not about trends, but rather about their own unique personal style, FA1 said they had to “crack down on personal aesthetics.” This crackdown involved implementing a rule that restricted the creation of article pages based on one’s own personal style. In this rule, FA1 clarified the goal of their community:
“The main goal is to document and be a reliable resource for people learning about different trends and communities in real life and online; multiple pages for things that don’t exist chips at the reliability.”
Reliability is an important feature of many wikis because people use them for acquiring information about a specific topic. Allowing the creation of articles that aren’t grounded in factual basis might jeopardize the goal of being a reliable source of information.
Community growth also tended to coincide with the addition of rules restricting spam. FV5 explained how in their Fediverse server for musicians in Boston, they added a rule to control spam by self-promoting local artists:
“I added a rule to make sure that if you’re promoting one event, just no more than two posts. I didn’t want the timeline to simply be a sort of clogged artery of promoters and no people actually talking to each other. Because I wanted it to be a space for people to talk to each other.”
Even though the problem was caused by good-faith local artists promoting their shows, FV5 felt that too many such posts threatened the community’s goal “to be a space for people to talk each other.” Similarly, R5’s community, a news discussion space for a region in Alabama, created a new rule to limit posts about topics that gain sudden virality and importance because they can set off a cascade of members posting threads about the same news article or topic. R5 explained their rationale for rulemaking to prevent such spam during the COVID-19 pandemic:
“Everyone was just losing their shit. Posting stuff every, I don’t know, 15-20 seconds at some point. I mean, it got ridiculous. So we set a rule in place that, okay, “Mega Thread Mode” is active. So what that essentially means is that there’s one thread at the top [of the subreddit’s webpage], it’s stickied. Anything related to that subject gets pulled into that thread.”
Implementing a rule restricting posts about suddenly important topics to a single “Mega Thread” helped to reduce the noise that was generated by the sudden influx of people posting the same news articles and allow space for other topics related to their community’s goals.
5.4.2. Problems from Decline
Leaders also attributed rule change to community decline manifesting in diminishing member activity or neglecting members’ problems. Some leaders responded to periods of diminishing member activity by changing the rules to reduce barriers to participation. In FA5’s community, members had complained on Talk Pages about “feeling kind of suffocated” by the rules’ requirements for contributing. The leaders listened and made these requirements less restrictive believing that they “need to not be so hardline on handling these things.” FV5 told us how they hadn’t changed rules yet but “might have to take rules away if they seem overly restrictive” for members of their Fediverse server. A more extreme example of changing the rules due to diminishing member activity comes from R3’s community, a place to discuss a popular US science-fiction TV series. Following the show’s conclusion, activity in the subreddit began to plummet. According to R3, their community’s moderators responded problem by evaluating not just the rules, but the community’s identity:
“We had a discussion about—okay, are we going to relax the rules? How far are we going to relax them? Are we going to allow people to post fan-fiction? Or even allow people to post memes?”
“Relaxing the rules” to allow fan-fiction and memes would open up new opportunities for members to make novel contributions to the community and sustain activity that isn’t solely rehashing discussions about the show’s content. That said, R3’s leaders also brought in to question what kind of community they would be: Would they preserve their exclusive focus on discussing the show’s content or become a more casual space with fan stories and low-effort memes?
Changing Leadership to Manage Decline
In rarer cases, communities recovered from neglecting members’ problems in the rules only by radically changing leadership first. For example, FA12 told us that the previous leadership team of their Fandom wiki was changed for being “lax with users who were being discriminatory, which was very disheartening to see for a while.” Harassment on wiki Talk Pages went unaddressed by moderators, leaving members feeling unsafe. The problem was eventually addressed through a “big overhaul of the moderator team” who also updated their rules to crack down on discriminatory language. In R9’s community, where members exchange plants through the postal mail, the previous leadership was “pushed out because they weren’t putting in any rules to protect people from scams.” As in FA12’s example, members did not feel safe and participation declined as the leaders failed to protect them. R9, along with some other core members, became their community’s new leaders through a community-wide vote and immediately added rules requiring verification photos, commenting on posts before privately messaging, and using PayPal Goods and Services for plant trades. Both of these cases show that communities may feel the need to completely change their leadership in order to survive periods of decline, especially if leaders overseeing the problems causing decline cannot or will not change the rules themselves.
5.5. Becoming Aware of the Need to Change the Rules
While problems themselves are primarily functional matters, we found that each stage in the processes of changing rules has significant relational aspects. The first stage of a rule change, however minor or major, is becoming aware of (new) problems so as to recognize the need to change, add, or delete as a rule. Here we elaborate two ways community leaders became aware of problems.
5.5.1. Reporting to Leadership
Community leaders often learned about problems not through direct experience, but from fellow leaders or community members. Leaders in FA1’s Fandom wiki, having recognized a problem, would “always discuss it with the [other] moderators,” to decide on whether they should implement a change. In many communities the leadership relied on members for awareness of problems. Rule change processes would happen “whenever the user base complains” (FA11), in which“someone has an issue with a rule, they bring it up—on the forum—to an admin” (FA3). Subreddits’ leaders learned about problems from members not only through direct messages or the ModMail system, but also in “meta” threads for conversations about the subreddit. Some subreddit leaders openly encouraged such feedback to learn how to improve governance. Still, some problems were acknowledged by leaders but left unaddressed. Leaders explained that their limited capacity as volunteers constrained their ability to find solutions to problems. For example, R1’s community leaders “never really had the time” to improve governance in their community because they were “kind of busy with real life.” It wasn’t until the COVID-19 pandemic that R1’s community leaders, either laid off from work or working from home, found extra time to spend on improving their community’s rules.
5.5.2. Reflecting After Enforcement
Our interviews highlighted how everyday practices of enforcing governance in the community could prompt reflection about rules after enforcement actions. For example, FA3 said that their community lacked rules in the early days and was like “the wild west.” The problems that arose for them eventually led them to craft rules that would address the behavior and content issues they had noticed. While FA3 points to an example where new rules were motivated by absence of rules, R8 noted how the enforcement of existing rules could reveal ways the current rules were insufficient. This would ultimately lead to discussions where community leaders took stock of their rule sets and reflected on how rules could be improved:
“At some point, we sit down and decide that the rules we have aren’t clear enough or that they don’t probably reflect the way that they have been enforced. And so we try to rewrite them all just based off of how we actually have been applying them.” (R8)
Our interviews suggest that these moments of reflection often involved formalizing and streamlining norms that emerged. FA4 explained that a reason for one of their biggest rule changes was “to overhaul” their communities’ rules to “properly consolidate all the unwritten rules” that they had accumulated. Leaders may also overhaul rule sets when they become too bloated with redundant or unnecessary rules that were also hard to enforce. FA12 explained they overhauled rules this way because their rule set was “really complicated and kind of difficult to keep track of.” In sum, sweeping rule changes usually occurred in mature communities where inaction would threaten community goals.
5.6. Deliberating on How to Solve Rule-Related Problems
When implementing a solution isn’t obvious or may disrupt community members’ experiences, leaders often engage in a deliberation process to better make sense of their current problem, what the current rules do to combat the problem, and how possible changes might ultimately resolve the problem. Interviews revealed how moderators and admins would create meeting spaces to discuss problems and how to handle them. Leaders on Fandom and Reddit often convened in group chats using software such as Discord where “everybody [is] adding their own [two] cents to it and then changing [rules] as a group.”(R7) FA1 described a situation where their team of moderators got on a voice call for over 3 hours reviewing and “cleaning” up their rule pages. FA4 similarly recalled a “lengthy process with the rest of the administrative team” when they drafted new rules for their Fandom wiki. Leaders described various procedures for these discussions. For example, R9 said that they want to make sure the other leaders “all get along well” so that they can “communicate disagreements and not have it be a big thing.” When disagreements did occur, the leaders were “respectful enough that they’ll get to a point where one of them is swayed one way or the other, or [the issue] is just tabled for another month or two.” Other procedures for rule change discussions involved teasing out potential issues. FA6 described this process:
“When we propose a rule or we’re thinking through it, I tried to test it for edge cases—like the gray areas—to see how we [handle] that or whether we have to change the rules or whether it fundamentally makes it not very enforceable.”
Leaders not only sought alignment on the problems and proposed rule solutions, but also sought to design rules that could feasibly be enforced by testing proposed rules against imagined “edge cases.”
While leaders discussed problems and solutions privately, they were also including community members to seek information and feedback. Importantly, members’ experiences mattered to leaders who didn’t want to implement rules that might disrupt their activities. FA6 called this “doing outreach” to understand what members wanted and to understand how they might react to a rule change. Discussions happened in spaces such as “Meta” threads on subreddits, and discussion forums or Talk Pages on Fandom. FA3 described this back-and-forth deliberation with members and leaders:
“We try to get input from other users, and if we come to a decision and everyone else is like ’No, that decision sucks’, then we’re like ’Well we think it works better because this, this, and this, but we see your point and we can compromise here, here, and here on that change.”
In this example, FA3’s community worked collaboratively on changing rules by finding compromises between the goals of the leaders and the experiences of community members. R4 explained that this approach to rule change involved “consultation with the community.” For example, community members would make problems visible to leaders by “flagging posts” and “complaining about this type of behavior.” Then, before implementing solutions, leaders would first communicate with members about their intended rule changes to ensure the changes wouldn’t disrupt their experiences in the community. Leaders on Reddit in particular would take into account the “upvote” feature that allows community members to vote on particular comments within meta discussion threads to help leaders understand “what people were wanting” (R5).
5.7. Making the Decision to Change the Rules
Finally, after encountering and observing problems and deliberating over potential solutions, community leaders needed to ultimately decide whether or not to change the rules. We consider this a separate step from deliberation because communities frequently differed in their decision-making procedures, which reflected differences in governance philosophies and division of labor. Governance procedures vary to a degree within each setting, but we observed some patterns: 1) autocratic decision-making for leaders of Fediverse communities, 2) oligarchic decision-making for leaders on Fandom, and 3) a (limited) implementation of democratic participation for rule change decision-making for some Reddit communities.
Simply put, the Fediverse leaders we spoke to rarely changed their rules. Most problems they encountered originated in external servers and were resolvable through the Fediverse’s “defederation” system of blocking entire servers from interacting with their communities. Internal problems were rare because new members were frequently vetted in the registration process to join. Leaders often felt like they deserved to govern their communities exactly as they desired, sanctioning rule breakers without worrying about backlash. FV4 described this approach to decision-making the Fediverse:
“Ultimately… I’m paying to run the server. I don’t really wish to pay to give someone who’s been flagrantly awful a platform. They can go build their own soapbox. I don’t wish to use my time and money and resources to give them a broader audience.”
FV4 highlights the fact that leaders of Fediverse communities not only have to volunteer their time to govern but also have to spend money to pay for hosting their Fediverse server—operating costs that leaders on platforms like Reddit and Fandom do not shoulder. Rather than join a community to challenge their rules, members are expected to know and respect “the rules and the norms of operation…for a community that you’re engaging in,” and treating joining a Fediverse community like you’re “engaging with a contract with a company” (FV13). As a result, the rules often reflected what leaders “want personally in a community experience or social media experience.” (FV1) When members are responsible for finding a Fediverse community that aligns with their beliefs, and leaders justify autocratic governance because they ”pay the bills,” the initial rules set by leaders often remain unchanged. In rare cases when rules were changed at all, they were mostly revised “to make them more clear” (FV11).
Decision-making processes were more involved for communities on Fandom and Reddit. In both platforms, communities were typically governed by more than one person. For Fandom communities, decision-making was typically “oligarchical”—decisions for a wiki community would be made within “a small collective of moderators and admins” (FA3). Leaders often sought consensus on how to handle problems and in discussions on changing the rules: “You always try to compromise or try to make your interests—make your points—to have these rules [or] not have these rules” (FA14). That said, we see hierarchical decision-making structures even within these small groups of leaders. FA1 pointed out that the original founder of their wiki community changed their rules with no deliberation from others when they came back from a hiatus “for half a year” because they “decided that the rules were too lax.” Fandom wikis’ hierarchies were often formalized in roles including bureaucrats (who held the most decision-making power), admins (who helped manage community-wide problems), and moderators (who did much of the rule enforcement). In FA11’s wiki, the “main bureaucrat” is “the final decision maker” who has “the central role of helping make the rules for the wiki and making the final decisions if there was a tie” when voting on rule changes.
While subreddit communities also made governance decisions within their small team of volunteer moderators, we heard of a few cases where leadership felt compelled to bring community members into the decision-making process. In the previous subsection we described the ways in which R4 believed “the good way to handle rule development is based off the community’s needs and [to] consult them the whole way.” R4 added rules “through a process of doing surveys every quarter” for their community in which they identified problems and asked members, “what do you think we can do to fix this?” Answers to the survey would be taken as suggestions to “build those rules” that would improve the community. One example they gave was how their community would handle advertising content from corporate blogs. R4’s team sent out a questionnaire giving members several options for how leadership handle advertisements in their community:
“We suggested a couple options that we’re like ‘Just remove them all’ or ‘Keep them but flair [^2] Flairing a post means that the title of the discussion post is labeled with a tag, in this case it would be “Advertisement.” them and mark them as advertising’, that kind of stuff. And that ended up being the result that we decided was probably best for the community. That’s forever another rule.”
The new rule that members helped decide on required “Educational articles or blogs that contain advertisements to a service are to flair their post with ‘Corporate Blog’.”
R9’s community had a history of neglect by former leaders who failed to implement rules that protect their members. When R9 and other leaders replaced the old guard, they decided to take an approach that incorporated community members into the decision-making process for changes. R9 explained that they do “community posts” once a month to allow members to give feedback about their experiences. When reviewing the feedback, if R9 sees “people a little bit cranky about some rules” that they have, they will “put out a poll…to see if they [members] want things to change or not.” Members vote on whether a change should be implemented, although sometimes, R9, told us, “many polls…end up basically being 50/50.” They shelve the problem and do another poll a month later. In this way, R9’s leaders ensure that they “change things as the community wants to be changed,” in order to make sure the community is going in the right direction, supporting community members’ goals, and safeguarding their experiences.
6. Discussion
Although online communities often adopt rules for the functional regulation of behavior or content, we find that this is only part of the story. Rules also reinforce a community’s distinctive identity within an online platform and establish a community’s legitimacy in broader community networks. This insight complements challenges prior social computing work from both the “social engineering” and “institutional design” perspectives and reveals underappreciated relational aspects of rulemaking. Some of these relational aspects are external, such as when rules draw distinctions or signal alignments between communities. Others are internal, as when rules reinforce identity or consensus among a community’s leadership team or between leaders and regular members. Both internal and external relationships are important in the processes of rulemaking; they draw attention to problems, help form solutions, and shape decisions about changing the rules—particularly as community leaders have limited capacity and time.
Across our three empirical settings, we observed system design shaping rulemaking processes by facilitating and constraining relations in different ways. For example, Reddit lets members use posts to voice, discuss, and vote on concerns that might warrant rule changes. This finding extends prior quantitative research in CSCW on changes in the rules of online communities (Frey and Sumner, 2019; Reddy and Chandrasekharan, 2023) and speaks to concerns that modern sociotechnical system designs restrict the capacity for members to affect change (Frey and Schneider, 2021; Schneider, 2022).
In the rest of our Discussion, we first consider the implications of the relational dynamics of online community rules in the design of governance systems for online communities. We then reflect on how our findings inform theories of institutional design and change in online communities as well as directions for future work in designing social computing systems.
6.1. Relational Dynamics of Online Community Rules
A prevailing assumption is that online community rules formalize social norms that regulate interactions within communities (Butler et al., 2008; Kraut and Resnick, 2012; Grimmelmann, 2017; Chandrasekharan et al., 2019; Seering, 2019). This assumption treats rules as instruments for directing human interactions in computer-mediated environments. We find that, in addition to advancing this functional purpose, also serve relational purposes as leaders design rules with respect to other communities, leaders, and users.
First, rules set and sustain community boundaries, particularly as a means of protecting community identity amid evolving membership dynamics. The challenge of protecting identity is important—particularly for communities prone to influxes of new members who may alter the community’s composition or undermine its focus (Lin et al., 2017). In response, communities discoverable on large platforms and open to broad participation often implement rules that restrict off-topic content in order to preserve their identity and focus. While prior work has suggested that rules contribute to identity formation (Butler et al., 2008; Fiesler et al., 2018; Seering et al., 2019), our findings show that communities also defend their identity over time by prescribing strict guidelines for participation, often through reference to related communities. We extend prior work by Hwang and Foote (2021) and TeBlunthuis et al. (2022) who emphasize the importance of membership size and niche audiences within ecosystems of interdependent communities. This relational use of rules helps explain the concept of “nested niches,” in which communities thrive through their embeddedness in a broader socio-informational ecosystem (Hwang and Foote, 2021). Rules provide tools for drawing and maintaining boundaries between specialist and generalist spaces—especially on platforms, such as Reddit, where communities often emerge from one another and share overlapping members and topics (Tan, 2018; TeBlunthuis and Hill, 2022).
{mdframed}
Design implication: Community leaders can improve compliance by conveying how the rules define their community identity and what makes them distinct from others (Matias, 2019b). Designers of social computing systems can support this by creating “introductory spaces”—e.g., side panels, about pages, pinned messages—that help newcomers understand what differentiates the community. Platforms might also provide leaders guidelines for customizing their rules within the platform’s ecosystem.
Second, rules signal legitimacy to potential newcomers and users—particularly leaders—of other communities. Doing so demonstrates commitment to governance as a means of attracting newcomers and establishing a good community reputation within a platform. This pattern was most salient on the Fediverse where reputation is vital for maintaining connections to other servers in the decentralized network environment. Many Fediverse leaders adopted the same rules as the most popular servers because these were a widely recognized standard for governance, enhancing the credibility of leaders who adopted them among their peers. A community’s rules thus signal to important external audiences that the leadership takes governance (i.e., content moderation) seriously and abides by inter-server social norms.
This finding relates to recent work on Fediverse rules proposing that the decentralized arrangement of Fediverse platforms leaves no higher authority to hold individual communities accountable, which may enable toxic and abusive online cultures (Nicholson et al., 2023). While Fediverse servers could adopt heterogeneous rules, isomorphic pressures from the surrounding environment (DiMaggio and Powell, 1983) push server leaders to adopt identical rules more than the other settings in our study.
Such institutional isomorphism is dramatic on the Fediverse, but reflects more general dynamics. Leaders of new online communities, facing uncertainty in how to succeed in their environment (Singh et al., 1986), are prone to failure (Kraut and Resnick, 2012) and may lack requisite resources or knowledge to write well-crafted rules on their own. Instead, they model their community’s institutions (including their rules) on older communities that they perceive to be legitimate or successful (DiMaggio and Powell, 1983). In our study, founding leaders believed that a robust rule set would both attract newcomers and strengthen relations with neighboring communities.
{mdframed}
Design Implication: Community leaders often seek legitimacy by aligning with broader, recognizable norms across the platform. Platforms can facilitate this by providing curated rule templates drawn from similar communities that prompt customization to establish a distinct identity.
Our findings also point to ways that sociotechnical system designs condition rulemaking. Rationales around legitimacy were paramount on the Fediverse but absent from interviews with communities on Fandom and Reddit. Meanwhile, rationales around identity were central on Reddit and Fandom but less salient in interviews with Fediverse leaders. We hypothesize that this difference is attributable to how communities on Reddit and Fandom are strongly held to platform-wide norms and policies, but Fediverse servers depend on connecting to a decentralized network of autonomous communities. Platform-wide norms reduce the uncertainty that creates isomorphic pressures, making the problem of identity differentiation more central. By contrast, because Fediverse communities engage in community-level blocking (Colglazier et al., 2024), it becomes much more important for them to signal legitimacy through their rules and behavior. Doing so indicates whether they are aligned with other communities and allows them to sustain a wide network. No similar affordance for one community to sanction another exists on Fandom and Reddit. As long as Fandom wikis and subreddits adhere to the platform-wide norms, governance signals do not affect their inclusion in the community ecology.
{mdframed}
Design Implication: Designers of decentralized networks or federated systems where community-level sanctions exist should create tools that help communities coordinate normative alignment, such as cross-community agreements like the Fediverse Friendly Moderation Covenant.
6.2. Navigating the Challenges of Rule Change
Recent work in social computing has emphasized the concern that users, despite being members of online communities, frequently lack effective mechanisms to voice concerns and implement changes that would improve their experiences  (Frey and Schneider, 2021; Schneider, 2022). Such deficiencies stem from sociotechnical system design where the single creator of an online community on most platforms is given autocratic control over all aspects of the community. Under this arrangement, change depends on the community creator, who often becomes a “benevolent dictator for life.” Our findings reveal some ways that community-driven change is possible under this arrangement, and describe how both leaders and members resolve community problems through rule changes.
First, additional problems may trigger a need to change the rules in an online community. Recent quantitative studies of rule changes in subreddits and Minecraft servers have found that increasing membership size and community age associate with leaders deciding to change the rules (Frey and Sumner, 2019; Reddy and Chandrasekharan, 2023). In contrast, we find that periods of decline, such as decreasing activity in the community, may also provoke rule changes. But instead of adding rules to restrict content, leaders may actually remove or relax rules to encourage wider participation.
{mdframed}
Design implication. Rule changes may stem from periods of participation growth as well as periods of decline. In times of flux, platforms should prompt leaders to re-evaluate their goals as well as their rules (Reddy and Chandrasekharan, 2023). Rapid growth may provoke rule expansion and clarification (Kiene et al., 2016), while rapid decline may signal a need for broader pathways for participation, such as allowing casual and off-topic contributions to niche discussion communities.
Next, we find that leaders, while knowing problems exist, may lack the capacity to implement effective rule changes to resolve those problems. March et al. (2000), analyzing the evolution of written rules at a university, argue that organizations are “thermostatic” in their attention to problems: as long as performance is satisfactory, problems may go unresolved or unnoticed. Online communities, we suggest, may be especially prone to neglect problems because their voluntary leaders have limited resources. Volunteer leaders of online communities do unpaid civic labor to sustain their communities (Dosono and Semaan, 2019; Matias, 2019a; Wohn, 2019) and often quit because of their lack of time brought about psychological distress (Schöpke-Gonzalez et al., 2024). Yet as communities mature, new problems often demand new rules (Butler et al., 2008; Kiene et al., 2016; Lin et al., 2017; Reddy and Chandrasekharan, 2023). Leaders we interviewed cited similar challenges in finding time to change their rules, carefully weighed the demands of potential new rule enforcement, and in one case, replaced the leadership team with members better positioned to improve community governance.
{mdframed}
Design implication. Volunteer leaders face time and attention constraints that impact their capacity for proactive governance. Platforms should provide tools that 1) help recruit and distribute lightweight governance tasks to trusted community members, and 2) surface and act on member feedback via structured feedback mechanisms like surveys or binding community votes (Frey and Schneider, 2021).
One way to reduce strain on leaders is to share governance work with the broader community. Doing so also taps into community members’ collective knowledge to help in identifying problems and co-creating solutions (Turco, 2016; Hwang et al., 2024). Through Frey and Schneider’s (Frey and Schneider, 2021) and Schneider’s (Schneider, 2022) participatory governance lens, none of the communities we investigated for this study encoded mechanisms for direct, democratic participation in governance; none had binding mechanisms through which community members’ voice directly influenced decision-making by leaders. Fediverse leaders we interviewed were particularly autocratic: they rarely changed their rules, never held discussions or encouraged feedback about the rules and governance, and made decisions with no input from members. In fact, Fediverse leaders often said they wanted to create their own server where they would get to decide how things should be. In contrast, Fandom leaders expressed concerns about managing delicate relationships with high quality wiki contributors and gathered their input before making decisions. Subreddit leaders seemed even more involved with and accountable to their members. Many reported rulemaking processes that involved much more participation from members and even “soft” binding mechanisms for member demands.
What makes Reddit so different? We propose that a combination of sociotechnical affordances and institutionalized norms sustain a stronger culture of participatory governance. On the affordances side, centralized discussion spaces and integrated voting tools help members collectively communicate their concerns and values to leaders, increasing pressure on them to solve problems. On the normative side, Reddit users have developed sitewide values emphasizing individual freedom and a deep skepticism of moderators who act unaccountably. While nothing binds subreddit moderators to accept member demands, soft binding mechanisms, along with low costs of exit, constrain moderators. For example, several subreddit leaders we interviewed were cautious about making decisions without input from their community members. They feared that if influential members’ concerns weren’t taken seriously, these members would take their expertise elsewhere. Similar examples with Fediverse leaders were non-existent in our sample.
Taken together, this arrangement may result in leader-member relationships that are mutually dependent: members need leaders to moderate content and enforce rules, while leaders need members to contribute content and sustain community activity levels. Evoking metaphors in content moderation from Seering et al. (2020), the leaders of these subreddits may be more like “governors” who satisfy the needs of their constituents and less like the “custodian” wiki leaders who curate and tidy wikis, or the Fediverse “dictators” who create and run servers however they see fit.
6.3. Limitations and Future Work
As with all interview-based studies, our findings are shaped by the scope of our sample, participants’ subjectivity and self-presentation, and the interpretative nature of our qualitative analysis. While our sampling strategy aimed to capture a diverse and meaningful range of communities across three different sociotechnical systems, our findings are not intended to be representative or generalizable. Instead, our findings instead provide in-depth insights into rulemaking practices among online community leaders. We only looked at Reddit, Fandom, and the Fediverse when we could have considered other platforms as empirical settings. On the other hand, very few platforms besides those three have publicly available methods of constructing datasets of community rules. Future work could extend or contribute to our findings by interviewing leaders from communities on other platforms. Because we exclusively interviewed leaders (i.e., admins and moderators), future work is needed to understand community members’ perspectives and experiences related to rules, especially with respect to how they select communities to join, their perceptions of legitimate community governance, and how they navigate the challenges of voicing problems in online communities to inspire change. We proposed that institutional isomorphism may be a mechanism guiding the behaviors of leaders to adopt similar rules, but longitudinal studies of rule diffusion across community networks should evaluate such a claim. Additionally, an ecological analysis of community rules could investigate whether and how rules are used to construct niches within platforms or more complex community ecosystems. Although recent work is beginning to analyze rule changes in online communities, we also propose further statistical work is needed to investigate the relationship between rule changes and member participation.
7. Conclusion
This study set out to understand the origins of rules in online communities. We conducted a comparative, grounded theory based analysis of 40 in-depth interviews with community leaders from subreddits, Fandom wikis, and Fediverse servers, sampled to ensure rule overlap and heterogeneity. We found that rules and rulemaking are instrumental for moderating content and behavior at the same time as they also allow community leaders to establish and manage community legitimacy and identity. These findings emphasize relational dimensions of online community governance consistent with sociological explanations of institutional isomorphism and ecological dynamics. Our findings also contrast with two prevailing approaches in social computing, both of which focus on functional aspects of rules and largely overlook where rules come from. We observed that rule creation and change practices are influenced by the sociotechnical design of the platform hosting the online community space, which may shape how problems are surfaced and who participates in problem-solving deliberation and decision-making. Rule change was often driven by problems emerging during community growth or decline, but hinged on the limited capacity of volunteer leaders, whether mechanisms for member feedback exist, and the relational dynamics between leaders and members. Together, our analysis reveals design opportunities to support effective online community governance given the relational origins of rules and rulemaking.
Acknowledgements.
Special thanks to Molly DeBlanc who participated in the early stages of this project with a focus on free/libre open source software (FLOSS) before her project split off into a separate study. This work was also conducted with a group of undergraduate research assistants who participated in data collection and preliminary analysis and who were instrumental to the project: Marlene Alanis, Paz Baum, Dylan Griffin, Amy Guo, Noah Hellyer, Aaron Phan, Eric Rosin, Caroline Rygg, Davida Yalley, Xiyu (Olivia) Yan, Yiqi (Grace) Zhu, Carolyn Zou, Emily Zou. This work was financially supported by the US National Science Foundation (grants IIS-2045055, IIS-1910202, IIS-2045055, DGE-1842165).
Appendices
Community Sampling Methods
To identify and select interview participants across Reddit, Fandom, and the Fediverse, we implemented a multi-step sampling strategy focused on communities with similar rules. This multi-step process involved 1) compiling datesets of rule sets for each of the three platforms, 2) conducting clustering analysis techniques based on rule text, and 3) manually analyzing communities within rule similarity clusters to select samples of community leaders for interview recruitment.
Data Collection
Three researchers on the research team—a postdoctoral fellow and two graduate students—constructed datasets capturing community rules on each platform.
  • Reddit: We crawled Reddit’s public API from 2020 to 2022 and recorded every publicly available subreddit’s rules data and community characteristics. We excluded subreddits with fewer than 1,000 members to ensure some expectation of activity, resulting in a final data set with  700,000 communities.
  • Fandom: We used a February 2020 snapshot from the Internet Archive [^3] https://archive.org/details/wikiteam_2020-02-09 that included the edit histories of  340,000 Fandom wikis. We excluded communities fewer than 200 unique article contributors and 10,000 daily article edits to ensure activity as well. We then downloaded policy and rule-related pages for each community.
  • Fediverse: We used an iterative sampling process beginning with the server mastodon.social (because of its centrality and status within the broader Fediverse network), discovering additional servers from its peer connections, and continuing discovering connections from connections. From this list of Fediverse servers, we then downloaded and extracted the rule texts from the /texttt/about/more HTML data on Mastodon servers and the /static/terms-of-service.html HTML data on Pleroma servers as a snapshot of data from September 25, 2021.
Cluster Analysis
To identify clusters of communities with similar rule language:
  • Reddit and Fandom: We concatenated the rule texts for each community to create a 2-dimensional encoding of their rules, reduced their dimensionality via the Doc2Vec algorithm using the gensim Python package and applied further dimensionality reduction using the UMAP Python package. We then used the HBDSCAN algorithm to group similar communities into clusters, giving us a list of clusters of communities with similar rule language.
  • Fediverse: For Fediverse communities, we generated a list of clusters of communities with similar rule language using the HBDSCAN algorithm after we computed pairwise n-gram similarity of communities’ rule texts using the following formula:
    $\frac{\text{count}(X_{1}\cap X_{2})}{\min\left(\text{count}(X_{1}),\text{count}(X_{2})\right)}$
Sample Selection
For two weeks in June, 2022, we assigned each empirical setting (Reddit, Fandom, and the Fediverse) to a subteam comprising of three undergraduate RAs and at least one graduate or post-doc researcher. To reduce the scope of recruiting efforts, these subteams evaluated the clusters to narrow down potential communities from which to sample leaders for interviews. Subteams analyzed clusters of varying sizes (e.g., small, medium, and large) based on:
  • Recent activity levels
  • Estimated community size
  • Thematic content
  • Rule characteristics
Subteams met weekly to discuss the results and created lists of communities to target for recruiting leaders for interviews.
Appendix A Rules in subreddits, Fandom wikis, and Fediverse servers
This section provides examples of rules from the communities from which we interviewed leaders. This first set of rules is from the subreddits r/cybersecurity (Figure 1). The second set of rules come from the Marvel Fandom wiki (Figure 2). Finally, the third example of rules comes from the SFBA.social Fediverse community (Figure 3). Communities on all three platforms tended to list additional detail about their rules, usually on other pages.
Figure 1. Rules for the r/cybersecurity subreddit community.
Figure 2. Rules for the Marvel Fandom wiki community.
Figure 3. Rules for the SFBA.social Fediverse community.
Appendix B Participating Community Characteristics
This section gives basic descriptions of the online communities we selected to sample interview participants. We give the community’s name, when it was created, an estimate of its member population size, and the number of rules the community had adopted at the time of data collection.
Community Created Size Rules
r/kdramarecommends 2014 68889 11
r/explainlikeimfive 2011 20945799 10
r/Supernatural 2008 202207 9
r/cybersecurity 2012 381366 9
r/huntsvillealabama 2008 60028 8
r/diabetes_t1 2013 41638 8
r/amcstocks 2021 67428 15
r/harrypotter 2008 1324162 14
r/takeaplantleaveaplant 2011 51156 15
Table 1. Reddit community characteristics. Size refers to the total number of subscribers to the subreddit.
Community Created Size Rules
Aesthetics Fandom 2018 4652 18
Marvel Fandom 2005 23481 41
Naruto Fanon Fandom 2007 4588 59
Valorant 2020 880 21
DC Wiki 2009 15025 33
Riordan Wiki 2007 13365 135
Wings of Fire Fanon 2013 10125 45
Brickpedia 2006 11926 28
Table 2. Fandom community characteristics. Size refers to the total number of users who have made at least 1 contribution to the wiki.
Community Created Size Rules Software
sfba.social 2019 3700 7 Mastodon
toot.wales 2018 981 3 Mastodon
medievalist.masto.host 2018 26 4 Mastodon
freeradical.zone 2017 278 15 Mastodon
bostonmusic.online 2020 2 6 Mastodon
neue.city Not Available 60 22 Pleroma (Akkoma)
hackers.town 2017 211 16 Mastodon
ausglam.space 2018 235 9 Mastodon
vocalounge.cafe 2020 49 6 Mastodon
mspsocial.net 2018 370 11 Mastodon
sonomu.club 2019 79 10 Mastodon
mastodon.radio 2018 1100 9 Mastodon
expired.mentality.rip Not Available Not Available Not Available Pleroma
vantil.refchat.net Not Available 31 Not Available Pleroma
mk.disqordia.space Not Available Not Available Not Available Misskey
glammr.us 2018 271 3 Mastodon
Table 3. Fediverse community characteristics. Size refers to the number of registered users on each server. Note: some servers’ community characteristics could not be found in any public online space.
Appendix C Interview Questionnaire
Questions about participation
  1. (1)
    What brought you to community <Community A>?
  2. (2)
    How did you choose to participate in <Community A>?
  3. (3)
    How long have you been involved with <Community A>?
  4. (4)
    How do you see your role within <Community A>?
  5. (5)
    How would you describe <Community A>?
  6. (6)
    Are you paid or otherwise compensated for your participation in <Community A>?
  7. (7)
    Do <Community A>’s choice of rules influence your decisions to choose to join or participate in <Community A> in any way? How so?
  8. (8)
    Have you ever had a negative experience with <Community A>? If so, was it related to the rules in any way?
Questions about rules in general
  1. (1)
    Can you tell me about <Community A>’s rules?
  2. (2)
    What do you believe is the purpose of rules in <Community A>?
  3. (3)
    Can you confirm that <Community A> uses the rules on their website?
  4. (4)
    Were you aware of <Community A>’s rules before you signed up?
  5. (5)
    Can you recall some specific rules of <Community A>? Why do these ones stick out to you?
  6. (6)
    How important are the rules in <Community A>?
  7. (7)
    What do you think would happen if you didn’t have <rule X>?
  8. (8)
    Are there any unwritten rules or norms in <Community A>?
  9. (9)
    Are there any rules that <Community A> has that you wish it didn’t? If so, why do you think they still have this rule?
  10. (10)
    How do newcomers learn the rules of <Community A>? How did you learn of the rules of <Community A> when you first joined?
  11. (11)
    Can you walk me through the first time you interacted with <Community A>’s rules / how you became familiar with <Community A>’s rules?
  12. (12)
    What kinds of behaviors does <Community A> view negatively? Why?
  13. (13)
    What kinds of behaviors does <Community A> view positively? Why?
  14. (14)
    (In follow up to the two questions above) How do <Community A>’s rules address <Behavior Y>?
  15. (15)
    Are there any rules that you wish that <Community A> had? Why? If so, why do you think the community doesn’t have a rule like that?
  16. (16)
    How do <Community A>’s rules align with your personal values? How do these rules reflect your values or not?
  17. (17)
    What are your personal values in community participation?
  18. (18)
    Can you tell me about your personal approaches to community participation?
  19. (19)
    How do <Community A>’s rules shape your contributions to this community? For example, do the rules affect the types of content or work that you produce and/or share?
  20. (20)
    In what types of settings or activities are your rules applied? Do the same rules apply in the mailing list as in the chat room? Do the same rules apply if you meet up in-person or at a convention?
  21. (21)
    Do you have different rules for different mediums or contexts? (e.g., mailing lists, chat rooms, forums, in-person meetups)
  22. (22)
    Are there things that tend to be agreed upon within <Community A> that aren’t written in formal rules?
  23. (23)
    How do the rules help, hurt, or otherwise impact the community?
  24. (24)
    Can you describe any problems in <Community A>?
  25. (25)
    Do the rules attempt to deal with these problems in any way?
Questions about the origins of rules
  1. (1)
    How did <Community A> choose its rules?
  2. (2)
    Why did you choose these rules?
  3. (3)
    Have you ever had to change the rules?
  4. (4)
    How have the rules changed, if ever?
  5. (5)
    Are there any other communities that <Community A> looked to when you were choosing rules?
  6. (6)
    What is the most recent rule that <Community A> has adopted? Tell me about that?
  7. (7)
    How do you think <Community A> decided to phrase <Rule X>?
  8. (8)
    Are there any rules that <Community A> used to have, but doesn’t have anymore? If so, do you know why this change was made?
  9. (9)
    What is the process for adopting or changing rules in <Community A>?
  10. (10)
    Do you know who set the rules in <Community A>? Can you tell me more about that?
  11. (11)
    Where do [you think] these rules come from? (looking at/through the rule list together)
Questions for moderators, administrators, or other leadership positions
  1. (1)
    Is moderating <Community A> part of your job or otherwise a paid activity?
  2. (2)
    Why do you work as a moderator of <Community A>?
  3. (3)
    How did you become a moderator of <Community A>?
  4. (4)
    How would you describe your work as a moderator of <Community A>?
Questions about the enforcement of rules
  1. (1)
    What do you think <rule X> means in practice? How does that feel different to what the rule means in theory, if at all?
  2. (2)
    How many people are involved in rule enforcement? What roles do they hold within the community?
  3. (3)
    Do you feel as though the rules are fairly/equally enforced in <Community A>? (Follow up: why do you think that?)
  4. (4)
    Do you feel as though the rules are consistently enforced in <Community A>? (Follow up: why do you think that?)
  5. (5)
    Do you feel as though it is possible for rules to be effectively enforced in <Community A>? (Follow up: why do you think that?)
  6. (6)
    How do you think moderators/administrators of <Community A> use the rules?
  7. (7)
    How strictly do the moderators of <Community A> enforce the rules? Why? What makes you feel like it is strict or not?
  8. (8)
    Do ordinary members of the community help enforce the rules or norms? (e.g., ”call each other out”)? If so, can you share an example of a time that happened?
  9. (9)
    If not, why do you think not?
  10. (10)
    In what ways are rules or norms enforced in <Community A>? Can you share an example of a time that happened?
  11. (11)
    Are there particular people who take responsibility for enforcing the rules or norms? Are these people normally in formal roles, like moderators or administrators?
  12. (12)
    What happens when there is conflict over how rules are enforced (internal to rules enforcement team)?
  13. (13)
    What happens when there is conflict over how rules are enforced (external to rules enforcement team/community wide)?
  14. (14)
    Are there any rules for moderators specifically?
Questions about related communities
  1. (1)
    Are you a part of other online communities?
  2. (2)
    How do the rules of <Community A> compare to the rules of related communities?
  3. (3)
    Do all communities like yours have a rule like <rule X>?
  4. (4)
    Can you name some other communities with rules?
  5. (5)
    Can you compare the rules of <Community A> to <Community B>?
  6. (6)
    How do the problems of <Community A> compare to those of <Community B>?
  7. (7)
    Do you think differences in rules shape the relationship between <Community A> and <Community B> in any way?
Questions about rules and platforms
  1. (1)
    Do you feel like you need to have certain rules because you are on <Platform A>?
  2. (2)
    How important are the rules on <Platform A>?
  3. (3)
    Can you tell me about the [most important] rules on <Platform A>? How does that translate into <Community A>?
  4. (4)
    How do <Community A>’s rules align or not with <Platform A>’s rules?
  5. (5)
    How do you think about <Platform A>?
  6. (6)
    How does <Platform A> compare with other online community platforms?
  7. (7)
    Are there other communities on <Platform A> you engage with regularly?
  8. (8)
    Do you feel as though the platform rules protect you personally?
  9. (9)
    Who do your community rules protect?
Questions about Background Demographics
  1. (1)
    Do you identify as being a minority or otherwise marginalized within your culture?
  2. (2)
    Do you identify as a minority or otherwise marginalized within <Community A>?
  3. (3)
    How does your identity or background shape your experience with <Community A>?
  4. (4)
    Do you feel like others who share your identity or background would feel comfortable participating in <Community A>? Why or why not?
  5. (5)
    Would you recommend <Community A> to people with minority or otherwise marginalized identities? Why or why not?
  6. (6)
    Can you describe any ways that the rules or norms of <Community A> could do more to help those with your identity or background?
  7. (7)
    Can you describe any ways that the rules or norms of <Community A> already help those with your identity or background?
  8. (8)
    Have you ever had a negative experience with <Community A> related to your identity or background?